Josh can but 10 roses for $8.00, how many can he buy with $18.00?

Answer by stanbon(75887)

Josh can buy 10 roses for $8.00, how many can he buy with $18.00?
------
each rose costs 800/10 = 80 cents.
---
# of roses = 1800/80 = 22.5
=============================
Cheers,
Stan H.